OpenStudy (anonymous):

the area of a circle is found by the equation a=pi r ^2 . if the area a of a certain circle is 16 pi square inches, find its radius r a. 8 pi inch b. 4 inch c. 32 inch d. 8 inch

myininaya (myininaya):

a=pi*r^2 16=pi*r^2 16/pi=r^2 (16/pi)^(1/2)=r 4/(pi)^(1/2)=r

myininaya (myininaya):

oh the area is 16pi

myininaya (myininaya):

ok so we have 16pi=pi*r^2

myininaya (myininaya):

divide both sides by pi so we have 16=r^2 what is the square root of 16?

myininaya (myininaya):

that will be your radius once you answer that question

OpenStudy (anonymous):

4

OpenStudy (anonymous):

thanks